OpenStudy (anonymous):

Use the Gauss method to find the sum of the first 12 terms in the pattern below. 8, 11, 14, 17, ...

OpenStudy (anonymous):

Can you tell me how much each number goes up by?

OpenStudy (anonymous):

Silent type, no problem. Each number goes up by 3.

OpenStudy (anonymous):

So starting from 17 you add 3 twelve times.

OpenStudy (anonymous):

I mean 8 times.

OpenStudy (anonymous):

17+3 = 20 (5) 20+3 = 23 (6) 23+3 = 26 (7) 26+3 = 29 (8) 29+3 = 31 (9) 31+3 = 34 (10) 34+3 = 37 (11) 37+3 = 40 (12)

OpenStudy (anonymous):

Thank you! (: x

OpenStudy (anonymous):

No problem :)